OpenStudy (anonymous):

find the unit rate in feet per second 300yards : 1 minute?

OpenStudy (cwrw238):

300 yds = 900 feet 900 ft are run in 60 secs unit rate = ft / sec = 900 / 60

OpenStudy (anonymous):

i don't get what you did could you explain please like by what do you time it? \[\frac{ 300 yards }{ 1 minute }\]

OpenStudy (cwrw238):

ok 1 yard = 3 feet so 300 yards = 3*300 = 900 ft 1 minute = 60 seconds unit rate = distance / time =900 / 60 feet per second

OpenStudy (anonymous):

so the unit rate would be\[\frac{ 900 }{ 60}\] what do the slash signs mean? could u put them in fractions pls (: im confused

OpenStudy (cwrw238):

'/' means divided by 900/60 = 900 --- 60

OpenStudy (anonymous):

i see :D so i divide that and the answer would be 15?

OpenStudy (cwrw238):

you can simplify this